OpenStudy (help!!!!):

11. Jim has 5 pieces of string. He needs to choose the piece that will be able to go around his 36-inch waist. His belt broke, and his pants are falling down. The piece needs to be at least 4 inches longer than his waist so he can tie a knot in it, but it cannot be more that 6 inches longer so that the ends will not show from under his shirt. Which of the following pieces of string will work the best? A. 3 feet B. 3 ¾ feet C. 3 ½ feet D. 3 ¼ feet E. 2 ½ feet

OpenStudy (help!!!!):

can you show me the equation?

OpenStudy (anonymous):

D too short which leaves C which is a the extreme length of 6 inches A. 3 feet too short B. 3 ¾ feet 3/4=9/12 too short C. 3 ½ feet 1/2=6/12 ok D. 3 ¼ feet 1/4=3/12 tooo short E. 2 ½ feet too short

OpenStudy (anonymous):

Well, a foot is 12 inches, so you wanna convert them all to inches.